To estimate the percentage of Binghamton residents who use public transportation, a simple random sample of 200 residents is taken. 75 of the people in the sample said they use public transportation, while the rest said they do not use public transportation.

Find the SE for the percentage of Binghamton residents who use public transportation in our sample.
Answer #1

Solution

Given that,

p = x/n=75/200=0.375

1 - p = 1-0.375=0.625

n = 200

SE=

=  [p ( 1 - p ) / n] =   [(0.375*0.625) / 200 ] = 0.03423